OpenStudy (anonymous):

Find the limit. Use l'Hospital's Rule if appropriate. If there is a more elementary method, consider using it. lim x^(9/x) x→∞ Please explain how to get to the answer.

OpenStudy (xapproachesinfinity):

L'Hopital can be used since it is \[\infty^0\]

OpenStudy (xapproachesinfinity):

you can use elementary method! do change of base of using natural exponential

OpenStudy (xapproachesinfinity):

@tuna.eng

OpenStudy (anonymous):

is it just ∞?

OpenStudy (xapproachesinfinity):

you can do \[\lim _{x\to \infty}x^{\frac{9}{x}}=\lim_{x\to \infty}e^{9\frac{\ln x}{x}}\]

OpenStudy (xapproachesinfinity):

lnx/x you can see that x grows faster than lnx so the limit will tend to zero

OpenStudy (anonymous):

I entered 0, it was wrong

OpenStudy (xapproachesinfinity):

no mu friend only lnx/x tends to zero not the whole thing

OpenStudy (xapproachesinfinity):

see if you can use L'hopital's rule with what i just did

OpenStudy (anonymous):

so if I plug in infinity into that...I'm not sure what would happen after

OpenStudy (xapproachesinfinity):

sorry i skipped some steps, i took into account that you are familiar with such operations with exponential and limits

OpenStudy (anonymous):

That's what's I got in my work also but that's where I am stuck

OpenStudy (perl):

(Just for clarity) $$ \Large { \lim _{x\to \infty}x^{\frac{9}{x}}= \lim _{x\to \infty} e^{\ln(x^{\frac{9}{x}})} =\lim_{x\to \infty}e^{\frac{9}{x}\cdot \ln (x)}\\~\\ =\lim_{x\to \infty}e^{\frac{9 \cdot \ln x}{x}} =e^{\lim_{x\to \infty} \frac{9 \cdot \ln x}{x}} } $$

OpenStudy (perl):

now if you plug in infinity, you get an indeterminate \( \large \bf {\frac{\infty }{\infty }} \), so you can use Lhopitals rule.

OpenStudy (xapproachesinfinity):

analyzing the behavior of lnx/x wouldn't be elementary method will it @perl ?

OpenStudy (anonymous):

I got 9/x

OpenStudy (xapproachesinfinity):

@tuna.eng take the limit of that as x goes to oo

OpenStudy (perl):

yeah, im not sure what they mean by 'more elementary method' . i guess easier method

OpenStudy (anonymous):

so 9/infinity... e^0=1?

OpenStudy (xapproachesinfinity):

hmm yes that's good

OpenStudy (perl):

$$ \Large { \lim _{x\to \infty}x^{\frac{9}{x}}= \lim _{x\to \infty} e^{\ln(x^{\frac{9}{x}})} =\lim_{x\to \infty}e^{\frac{9}{x}\cdot \ln (x)}\\~\\ =\lim_{x\to \infty}e^{\frac{9 \cdot \ln x}{x}} =e^{\lim_{x\to \infty} \frac{9 \cdot \ln x}{x}} \\ ~\\ \Rightarrow e^{\lim_{x\to \infty} \frac{9 }{x}} =e^{ \frac{9 }{\infty }} = e^0 = 1 } $$

OpenStudy (anonymous):

okay! I understand ! Thank you!!!!!!!

OpenStudy (perl):

$$ \Large { \lim _{x\to \infty}x^{\frac{9}{x}} \equiv \lim _{x\to \infty} e^{\ln(x^{\frac{9}{x}})} \equiv\lim_{x\to \infty}e^{\frac{9}{x}\cdot \ln (x)}\\~\\ \equiv\lim_{x\to \infty}e^{\frac{9 \cdot \ln x}{x}} \equiv e^{\lim_{x\to \infty} \frac{9 \cdot \ln x}{x}} \\ ~\\ \Rightarrow e^{\lim_{x\to \infty} \frac{9 }{x}} \equiv e^{ \frac{9 }{\infty }} \equiv e^0 \equiv 1 } $$

OpenStudy (perl):

but thats being pedantic :)

OpenStudy (perl):

= sign does not always mean identity, it can be a conditional equation

OpenStudy (xapproachesinfinity):

hmm true :)

OpenStudy (perl):

or you could say, the equalities above are true for all x in the domain, then you dont need to use identity

OpenStudy (jhannybean):

Very nicely written @perl :)

OpenStudy (perl):

thanks

OpenStudy (anonymous):

how did you know how to use e?

OpenStudy (xapproachesinfinity):

i often use equalities not paying attention to domain and the condition only very few cases that i consider the domain lol (when i need it)

OpenStudy (perl):

this is similar to $$ \Large{ \lim _{x \to 0} ~x^x \\~\\\\~~\large \text{or}\\ \lim _{x \to \infty} x^\frac{1}{x} } $$

OpenStudy (jhannybean):

Just by looking at the problem, you see a variable as a fraction in the exponent. I think to simplify and solve for it (taking the limit of it) we either would use the log function, or the natural function the "bring down" the fraction in order to simplify it.

OpenStudy (jhannybean):

That's my quick assessment of this problem. Not 100% correct I don't think haha.

OpenStudy (xapproachesinfinity):

this function \[x^x\] does it have applications i don't see it that often.

OpenStudy (perl):

its only defined on reals for x>0 , and it grows superfast,

OpenStudy (anonymous):

okay jhannybean...that makes sense. Thank you all!

OpenStudy (xapproachesinfinity):

hmm yes, seems only x>0 works

OpenStudy (jhannybean):

So @perl, I wanted to why it was \[\huge e^{\ln(x^{9/x})}\] instead of \[\huge e^{\log_e(x^{9/x})}\]

OpenStudy (perl):

$$ \Large { \lim_{x \rightarrow \infty} e^{f(x)} \neq lim_{x \rightarrow \infty}~ f(x) } $$

OpenStudy (xapproachesinfinity):

alright i turned it to more than one question here lol! i hope i'm not confusing the poster

OpenStudy (jhannybean):

I think you misunderstood my question.

OpenStudy (perl):

oh, ln is shorthand for log_e

OpenStudy (xapproachesinfinity):

yeah no difference, there are the same just terminology

OpenStudy (jhannybean):

\(\color{blue}{\text{Originally Posted by}}\) @Jhannybean \[\huge \color{red}{e}^{\ln(x^{9/x})}\] instead of \[\huge \color{red}{e}^{\log_e(x^{9/x})}\] \(\color{blue}{\text{End of Quote}}\) And why are we using this??

OpenStudy (perl):

$$ \Large \ln(x) \equiv \log_e (x) $$

OpenStudy (jhannybean):

the highlighted red portion.

OpenStudy (xapproachesinfinity):

to change to base which is more understandable to us

OpenStudy (jhannybean):

Ok.

OpenStudy (perl):

this allows us to utilize the properties of logs, one of them is that we can bring down exponent in front

OpenStudy (xapproachesinfinity):

e and ln are always the choice when dealing with such situation there are natural and we understand them better than other bases

OpenStudy (xapproachesinfinity):

eh of course what perl also said about using the power property of logs

OpenStudy (jhannybean):

Got it \(\checkmark\) I just think of \(\large e^{\color{red}{\ln}(f(x))}\) the red portion as part of a packaged deal. If you're using e, then either use \(\ln\) or \(\log_e\) with it.

OpenStudy (perl):

in general $$ \Large { \lim_{x \rightarrow a} f(x)^{g(x)} =\lim_{x \rightarrow a}~ e^{\ln (f(x)^{g(x)})} = \lim_{x \rightarrow a}~ e^{g(x) \cdot \ln (f(x))} } $$

OpenStudy (jhannybean):

Oh I was about to write that, haha.

OpenStudy (perl):

errr

OpenStudy (xapproachesinfinity):

hmm just a curiosity question whenever we deal with such situation you just wrote we mostly assume f and g are continuous would there be a problem when one is not continuous

OpenStudy (xapproachesinfinity):

hahaha :)

OpenStudy (jhannybean):

If one isn't continuous, how would you be able to find its limit?

OpenStudy (jhannybean):

I thought finding the limit, or calculating the derivative of a function has to work if the function is continuous.

OpenStudy (xapproachesinfinity):

the limit might exist even if it is not continuous

OpenStudy (perl):

notice that in the problem I brought the limit inside the exponential function. $$ \Large \lim_{x \to a } ~e^{f(x)} = e^{\lim_{x \to a } f(x)} $$ this is valid because e^x is continuous function

OpenStudy (xapproachesinfinity):

yea i got that part exp is a continuous function what if f and g have discontinuity would that change the way we deal with the limit?

OpenStudy (perl):

this is a valid step*

OpenStudy (perl):

as long as it is continuous near the value x = a , or x = infinity

OpenStudy (perl):

so for instance lim x^x , as x ->0 , there is a discontinuity at x = 0

OpenStudy (perl):

so technically you really want $$ \Large{ \lim _{x \to 0^{+}} ~x^x } $$

OpenStudy (xapproachesinfinity):

i see, makes sense thanks :)

OpenStudy (perl):

the two valued limit doesn't really make sense. since x is not defined for x<0

OpenStudy (perl):

but you could define x for complex numbers and save the two valued limit or something

OpenStudy (xapproachesinfinity):

hmm going to complex analysis! i see

OpenStudy (perl):

you can say , find the limit of z^z as z->0 , the real part http://www.wolframalpha.com/input/?i=graph+x^x

OpenStudy (perl):

(-.1) ^(-.1) is not a real number , though

OpenStudy (perl):

Re( (-.001)^(-.001)) ~ 1

OpenStudy (xapproachesinfinity):

hmm, still not acquainted with complex analysis hehe

OpenStudy (perl):

if you have a TI 83

OpenStudy (perl):

(-.001)^(-.001) = 1.0069267 - i * .0031633 (approximately) So the 'real' part is going to 1 . but the 'imaginary' part is going to zero

OpenStudy (perl):

that step was Lhopitals rule

OpenStudy (xapproachesinfinity):

i will try it in ti

OpenStudy (perl):

well that step wasn't obvious :)

OpenStudy (perl):

$$ \Large { \lim _{x\to \infty}x^{\frac{9}{x}}= \lim _{x\to \infty} e^{\ln(x^{\frac{9}{x}})} =\lim_{x\to \infty}e^{\frac{9}{x}\cdot \ln (x)}\\~\\ =\lim_{x\to \infty}e^{\frac{9 \cdot \ln x}{x}} =e^{\lim_{x\to \infty} \frac{9 \cdot \ln x}{x}} \\ ~\\ \Rightarrow^{L'hopitals}\\ e^{\lim_{x\to \infty} \frac{9 }{x}} =e^{ \frac{9 }{\infty }} = e^0 = 1 } $$

OpenStudy (xapproachesinfinity):

@jhannybean don't use parenthesis for latex use square brackets

OpenStudy (perl):

what do you mean xapproachesinfinity

OpenStudy (jhannybean):

Why?

OpenStudy (xapproachesinfinity):

for the limits and other stuff to appear nice like perl's lol

OpenStudy (xapproachesinfinity):

oh no sorry for the confusion :)

OpenStudy (perl):

$$ \Huge \cancel{e^{ln(x)}} = x $$

OpenStudy (perl):

i want to cancel only the e and ln

OpenStudy (perl):

$$\Huge \color{red}{\cancel{\color{black}{e^{ln}}}}{}^{(x)} = x$$

OpenStudy (jhannybean):

\[\huge x^{\frac{3}{\ln(5x^3)}}\] and how about this one?

OpenStudy (perl):

where is the limit going to

OpenStudy (jhannybean):

Thought \(-\infty\) was possible, but it's not, so just \(+\infty\)

OpenStudy (xapproachesinfinity):

if you want use l'hopitals rule make x goes to 1 plus

OpenStudy (xapproachesinfinity):

oh no forget the factor there

OpenStudy (perl):

if we plug in infinity we get infinity ^ 0

OpenStudy (perl):

since ln(infinity) = infinity 1/ ln(infinity ) = 0

OpenStudy (xapproachesinfinity):

yes! i disregarded infinity and looked at other cases